flourine gas exerts pressure of 297 torr when the volume changes to 250 mL at STP.what was the original volume?


1
Expert's answer
2021-05-21T02:34:00-0400

Given:

Initial Pressure P1=297torr

Let initial volume be V1

According to the question ,Final Volume =250mL

In final condition gas is at STP. it means T=273K,and P2=760torr


So, at constant temperature :-

P1V1=P2V2(Boyle's Law)

297×V1=760×250


V1="\frac{760×250}{297}=639.73mL"


So, Original Volume ="639.73mL"